Two spherical drops of mercury each have a charge of 9.0×10−2 nC and a potential of 310 V at the surface. The two drops merge to form a single drop.

What is the potential at the surface of the new drop?

Express your answer to two significant figures and include the appropriate units
Answer #1

Electric potential due to point charge is given by

V=KQ/r

Therefore

=>r=KQ/V =(9*109)(9*10-2*10-9)/310

r=2.613*10-3 m

After two drops merge into single drop ,

V' =2V

(4pi/3)r'2=2(4pi/3)r3

r' =(2)1/3(2.613*10-3)=3.292*10-3 m

Total charge

Q'=2Q =2*(9*10-11)=1.8*10-10 C

Therefore potential at the surface of the new drop is

V'=KQ'/r'=(9*109)(1.8*10-10)/(3.292*10-3)

V' = 492.1 Volts =4.9*102 Volts (approx)